# Do Zombies Understand? A Choose-Your-Own-Adventure Exploration of Machine Cognition

### v1

@@ 1 Introduction: A Thought Experiment
To make this concrete, consider the following thought experiment: you are presented with $Z$, a new newfangled chatbot. $Z$ is implemented in computer hardware and performs only mathematical manipulations of its input. It is completely open-source — you have access to its code, training data, weights, hyperparmeters, and any other implementation detail. You interact with $Z$ and discover that it excels on all NLP benchmarks, and will do so on any possible test you will come up with in the future. In essence, $Z$ is the chatbot equivalent of the philosophical zombie (Kirk, 1974; Chalmers, 1996); it outperforms humans on all tasks, supposedly without having subjective experience. Do you consider $Z$ as capable ofunderstanding”?
@@
If you feel uncomfortable with either of these options turn to Section 5, where we address potential objections to our setup and assumptions.
@@ 3 Zombies do Understand:
Definition 1: Functional Understanding
A model $Z$ functionally understands a task $T$ if its performance on $T$ is as good (or better than) a human who is an expert in $T$.
This framing helps explain the common practice for testing understanding in models through longstanding challenges, such as chess, Go, or language generation. McCarthy (1990) figuratively called such tasks the Drosophila of AI, drawing a parallel between research in AI and biology, where model organisms (e.g., the Drosophila fly) are chosen for wide benchmark experimentation with findings generalizing beyond that specific organism.
@@
Perhaps the most well-known examples are the games of chess and Go. Chess served as a proxy task for understanding for nearly 50 years. Early works, such as Shannon (1950) and Turing (1953), already deemed chess a benchmark for machine intelligence. With the advent of deep learning models, chess engines now vastly outplay any human opponent (Silver et al., 2017). For all intents and purposes, these models functionally understand chess according to Definition 3. Consequently, chess was abandoned as a useful benchmark for understanding. [^1] New chess engines are still being developed, albeit without any claims about general understanding beyond chess. Instead, the game of Go was adopted as a marker for understanding Bouzy and Cazenave (2001); Van Der Werf (2004), until here, too, models have outplayed even the best human players Silver et al. (2016).
@@ 4 Zombies don’t Understand: Consciousness is a Prerequisite for Understanding
Definition 4 formulates this notation with regards to a model $M$ and a task $T$:
Definition 2: Conscious Understanding $M$ consciously understands $T$ if both hold: (1) $M$ functionally understands $T$ (§Def. 3). (2) $M$ is conscious – it has immediate subjective experience. In Nagel (1974)’s words there is something that it is like to be $M$.
As we highlight below, this notion of understanding has been articulated by seminal works in the field of AI and NLP. In a section titled Argument from Consciousness from his famous paper, Turing (1950) cites (Jefferson, 1949): [^2] Emphasis is our own in all quotes.
@@
To move forward on conscious understanding as articulated in Definition 4, we suggest following literature in neuroscience regarding the neural-correlate-of-consciousness (NCC; Koch et al., 2016). This field is dedicated to recognizing the neural dynamics in biological organisms that are associated with consciousness experience. For example, the Integrated Information Theory (IIT; Tononi et al., 2016), and more specifically the weak IIT, links elements of consciousness with wider information flow metrics, like recurrent processing, or global workspace. These findings can inform cognitively-inspired architectures, e.g., spiking neural networks (Mediano et al., 2022).
@@ 5 Other Possible Answers
Whether $Z$ understands depends on its implementation (training data, architecture, hyperparameters, etc.), but that has nothing to do with conscious experience.
This argument assumes that there may exist implementations of $Z$ which will show that it indeed understands, e.g., if they involve complex feature manipulation or explicit reasoning steps, while there may exist other implementations which imply that $Z$ does not understand, e.g., if all $Z$ does is leverage spurious correlations or memorize an immense look up table, similar to the Chinese room argument (Searle, 1980). We posit that this argument is compatible within our discussion around functional understanding (§3). For example, if $Z$ leverages spurious correlations, then by definition there are samples which do not exhibit these correlations and which will stump $Z$ (otherwise they would not be spurious), contradicting our assumption that $Z$ is a philosophical zombie, and does not make non-human errors. Similarly, since human language can produce an infinite amount of meaningful texts (Chomsky, 2002), and $Z$ can only memorize a finite amount of samples (as it is implemented in finite hardware), then there must be samples outside of its memory on which it is bound to fail. This again contradicts our initial assumption that $Z$ does not fail where humans do not fail.
@@
We do not aim to define apriori what constitutes understanding, and do not argue that there is a single “correct” definition. Instead, we try to tease apart what researchers mean when they use the term, specifically highlighting the role that consciousness plays in it, and examine how AI research may be explained through this lens. In fact, we claim that answering the question elucidates different definitions for understanding (Definitions 3,4). We invite researchers to engage with this question to examine their definition for understanding.

### v2

@@ 1 Introduction: A Thought Experiment
To make this concrete, consider the following thought experiment: you are presented with $Z$, a newfangled chatbot. $Z$ is implemented in computer hardware and performs only mathematical manipulations of its input. It is completely open-source — you have access to its code, training data, weights, hyperparmeters, and any other implementation detail. You interact with $Z$ and discover that it excels on all NLP benchmarks, and will do so on any possible test you will come up with in the future. In essence, $Z$ is the chatbot equivalent of the philosophical zombie (Kirk, 1974; Chalmers, 1996); it outperforms humans on all tasks, supposedly without having subjective experience. Do you consider $Z$ as capable ofunderstanding”?
@@
If you feel uncomfortable with either of these options, turn to Section 5, where we address potential objections to our setup and assumptions.
@@ 3 Zombies do Understand:
Definition 1
Functional Understanding. A model $Z$ functionally understands a task $T$ if its performance on $T$ is as good (or better than) a human who is an expert in $T$.
This approach to understanding is articulated in Dummett (1996)’s discussion around intelligence: [^1] Emphasis is our own in all quotes.
If a Martian could learn to speak a human language, or a robot be devised to behave in just the ways that are essential to a language speaker, an implicit knowledge of the correct theory of meaning for the language could be attributed to the Martian or the robot with as much right as to a human speaker, even though their internal mechanisms were entirely different.
Dummett (1996)
This framing helps explain the common practice for testing understanding in models through longstanding challenges, such as chess, Go, or language generation, or in many NLP benchmarks, such as text comprehension (Wang et al., 2018; Hendrycks et al., 2021; bench authors, 2023; Liang et al., 2023) or formal semantic representation (Oepen et al., 2014; Nivre et al., 2016).
McCarthy (1990) figuratively called such tasks the Drosophila of AI, drawing a parallel between research in AI and biology, where model organisms (e.g., the Drosophila fly) are chosen for wide benchmark experimentation with findings generalizing beyond that specific organism.
@@
Perhaps the most well-known examples are the games of chess and Go. Chess served as a proxy task for understanding for nearly 50 years. Early works, such as Shannon (1950) and Turing (1953), already deemed chess a benchmark for machine intelligence. With the advent of deep learning models, chess engines now vastly outplay any human opponent (Silver et al., 2017). For all intents and purposes, these models functionally understand chess according to Definition 1. Consequently, chess was abandoned as a useful benchmark for understanding. [^2] New chess engines are still being developed, albeit without any claims about general understanding beyond chess. Instead, Go was adopted as a marker for understanding Bouzy and Cazenave (2001); Van Der Werf (2004), until Go models outplayed the best human players Silver et al. (2016).
@@ 4 Zombies don’t Understand: Consciousness is a Prerequisite for Understanding
This notion is formulated with regards to a model $M$ and a task $T$ in the following definition:
Definition 2
Conscious Understanding. $M$ consciously understands $T$ if both hold:
  1. 1.
    $M$ functionally understands $T$ (§Def. 1).
  2. 2.
    $M$ is conscious – it has immediate subjective experience. In Nagel (1974)’s words there is something that “it is like” to be $M$.
As we highlight below, this notion of understanding has been articulated by seminal works in the field of AI and NLP. In a section titled Argument from Consciousness from his famous paper, Turing (1950) cites (Jefferson, 1949):
@@
To move forward on conscious understanding as articulated in Definition 2, we suggest following literature in psychology and neuroscience regarding tests for consciousness (for review, see Bayne et al. (2024)) and specifically the neural-correlate-of-consciousness (NCC; Koch et al., 2016). This field is dedicated to recognizing the neural dynamics in biological organisms associated with consciousness experience. For example, the Integrated Information Theory (IIT; Tononi et al., 2016), specifically the weak IIT, links elements of consciousness with wider information flow metrics, like recurrent processing or global workspace. These findings can inform cognitively-inspired architectures, e.g., spiking neural networks (Mediano et al., 2022).
@@ 5 Other Possible Answers
Whether $Z$ understands depends on its implementation (training data, architecture, hyperparameters, etc.), but it has nothing to do with conscious experience.
This argument is in line with Block (1981)’s definition of Psychologism, which assumes that there may exist implementations of $Z$ which will show that it indeed understands, e.g., if they involve complex feature manipulations or explicit reasoning steps, while there may exist other implementations which imply that $Z$ does not understand, e.g., if all $Z$ does is leverage spurious correlations or memorize an immense look up table, similar to the Chinese room argument (Searle, 1980).
We argue that the concerns regarding specific implementations not being indicative of understanding can be mitigated with our requirement that $Z$ excels on all possible NLP benchmarks, while also being implemented on a physical hardware. For example, if $Z$ leverages spurious correlations, then by definition there are samples which do not exhibit these correlations and which will stump $Z$ (otherwise they would not be spurious), contradicting our assumption that $Z$ is a philosophical zombie, and does not make non-human errors. Similarly, since human language can produce an infinite amount of meaningful texts (Chomsky, 2002), and $Z$ can only memorize a finite amount of samples (as it is implemented in finite hardware), then there must be samples outside of its memory on which it is bound to fail. This again contradicts our assumption that $Z$ does not fail where humans do not fail.
@@
We do not aim to define apriori what constitutes understanding, and do not argue that there is a single “correct” definition. Instead, we try to tease apart what researchers mean when they use the term, specifically highlighting the role that consciousness plays in it, and examine how AI research may be explained through this lens. In fact, we claim that answering the question elucidates different definitions for understanding (Definitions 1,2). We invite researchers to engage with this question to examine their definition for understanding.
@@ 6 Discussion
Acknowledgments
We would like to thank the anonymous reviewers for their thoughtful comments and suggestions, and Dr. Anat Arzi, Dr. Yael Bitterman, and Prof. Oron Shagrir, for many insightful and productive discussions. This work was partially supported by a grant from the Israeli Ministry of Science and Technology (grant no. 2336).